OpenStudy (anonymous):

Find the exact real value of arccos (1/2)

hartnn (hartnn):

what do u think ?

OpenStudy (anonymous):

i don't even know what arccos is.........

hartnn (hartnn):

you know where cos is ? see the unit circle http://en.wikipedia.org/wiki/File:Unit_circle_angles_color.svg

hartnn (hartnn):

when is cos value = 1/2 ??? see the x co-ordinate

OpenStudy (anonymous):

i typed it into google and this is what i got... 1.04719755 rad

hartnn (hartnn):

yeah, that equals pi/3 too! see pi/3 angle , you'll see (1/2, sqrt 3/2) means cos pi/3 = 1/2 sin pi/3 = sqrt 3/2 :)

OpenStudy (anonymous):

thank you so much for all your helpppp (:

hartnn (hartnn):

welcome ^_^